# A Modular DTaaS Architecture for Predictive Slice Management in 6G Systems

### v1

@@ I Introduction
Among the key architectural enablers of 6G, network slicing has emerged as a crucial mechanism for supporting service diversity and isolation. A network slice is a logical end-to-end construct that allocates radio, computing, and storage resources to meet the specific quality-of-service (QoS) requirements of diverse service classes, including enhanced Mobile Broadband (eMBB), ultra-Reliable Low-Latency Communication (URLLC), and massive Machine-Type Communication (mMTC). This logical partitioning allows network operators to provide customized performance guarantees, maximize resource efficiency, and ensure that each service category receives the necessary quality [3].
@@ II Related Work
Network slicing is a foundational technology for service differentiation in 5G and 6G. In [4], a reinforcement learning (RL) approach is proposed for network slicing in 5G networks, enabling dynamic resource allocation across diverse services. While it enhances efficiency, the method is reactive and may incur SLA violations in highly dynamic 6G environments. Similarly, [5] introduces a hybrid deep learning model for enhanced network slicing in 6G, focusing on security, latency reduction, and throughput improvement. However, it lacks integration with real-time digital representations for proactive forecasting. Our DTaaS framework overcomes these by embedding per-slice digital twins (SDTs) with predictive models for proactive provisioning and SLA assurance.
Digital twins (DTs) offer promising capabilities for network optimization. In [6], a digital-twin-driven approach is presented for end-to-end network slicing toward 6G, integrating AI to enable slicing orchestration. The work demonstrates improved management but does not emphasize edge deployment or multi-domain telemetry for fine-grained predictions. Likewise, [7] proposes a modeling and deployment framework for digital twins in 5G networks, creating virtual replicas for monitoring and optimization. Although applicable to 6G, it focuses on physical layer aspects without addressing service-level slicing dynamics. Our DTaaS extends DTs to a modular, edge-embedded service, incorporating cross-layer telemetry for slice-specific predictive orchestration.
@@
[4] RL-based dynamic resource allocation for slicing Reactive, may cause SLA violations in dynamic 6G
[5] Hybrid DL for slicing with security and latency focus No integration with real-time digital twins
[6] DT-driven orchestration for end-to-end slicing Lacks edge deployment and multi-domain telemetry
[7] DT modeling for 5G virtual replicas Focus on physical layer, not slice-level orchestration
[8] Deep RL for online slice resource allocation Centralized, unsuitable for URLLC latency
[9] Deep RL for flexible beyond-5G slicing Static training, lacks continuous feedback
[10] Federated learning for slicing in vehicular networks No synchronized twin representation for full state tracking
[11] Zero-touch orchestration for large-scale slices Focus on policies, not predictive SLA management
Predictive orchestration is essential for adaptive 6G networks. In [8], deep reinforcement learning is applied for online resource allocation in network slicing, achieving efficient adaptation to varying demands. Its centralized nature, however, introduces latency issues unsuitable for URLLC slices. In [9], deep RL is used for network slicing in beyond-5G and 6G systems to flexibly accommodate wireless services. While scalable, it relies on static training without continuous feedback loops for SLA risk prediction. Our framework deploys lightweight SDTs at the edge, merging sequential models with closed-loop control to reduce latency and bolster SLA compliance.
AI-native paradigms for 6G orchestration have also gained attention. In [10], federated learning is employed for network slicing in 6G systems tailored to autonomous vehicles, improving connectivity through distributed training. The approach enhances privacy but lacks synchronized twin representations for holistic state tracking. Similarly, [11] proposes a framework for zero-touch management and orchestration of massive network slice deployments in 6G, supporting automated lifecycle handling. It prioritizes policy enforcement over predictive resource management. Our DTaaS distinguishes itself by integrating SDTs into the orchestration loop, facilitating proactive scaling and reconfiguration via real-time telemetry and analytics.
@@ V-A Simulation Environment
To evaluate the proposed DTaaS framework, a simulation environment was developed in Python using the SimPy discrete-event simulation library [13]. The testbed emulates a 6G slicing scenario with heterogeneous services, including eMBB, URLLC, and mMTC. Each service category is instantiated as an independent network slice with unique traffic characteristics, latency constraints, and reliability targets.
@@ V-A1 Baseline Methods
  • Centralized Deep Reinforcement Learning (C-DRL): Inspired by recent studies such as [8, 9], this baseline applies a centralized deep reinforcement learning model trained to allocate resources among active slices. It learns optimal allocation policies over time but lacks real-time predictive awareness and introduces control latency due to centralized inference. Full RL implementations in live networks are complex; therefore, this baseline simulates the logical behavior of centralized agents rather than training on live data, enabling fair comparison with DTaaS.

### v2

@@ I Introduction
Among the key architectural enablers of 6G, network slicing has emerged as a crucial mechanism for supporting service diversity and isolation. A network slice is a logical end-to-end construct that allocates radio, computing, and storage resources to meet the specific quality-of-service (QoS) requirements of diverse service classes, including enhanced Mobile Broadband (eMBB), ultra-Reliable Low-Latency Communication (URLLC), and massive Machine-Type Communication (mMTC). This logical partitioning allows network operators to provide customized performance guarantees, maximize resource efficiency, and ensure that each service category receives the necessary quality [BILEN201824].
@@ II Related Work
Network slicing is a foundational technology for service differentiation in 5G and 6G. In [ref1], a reinforcement learning (RL) approach is proposed for network slicing in 5G networks, enabling dynamic resource allocation across diverse services. While it enhances efficiency, the method is reactive and may incur SLA violations in highly dynamic 6G environments. Similarly, [ref2] introduces a hybrid deep learning model for enhanced network slicing in 6G, focusing on security, latency reduction, and throughput improvement. However, it lacks integration with real-time digital representations for proactive forecasting. Our DTaaS framework overcomes these by embedding per-slice digital twins (SDTs) with predictive models for proactive provisioning and SLA assurance.
Digital twins (DTs) offer promising capabilities for network optimization. In [ref3], a digital-twin-driven approach is presented for end-to-end network slicing toward 6G, integrating AI to enable slicing orchestration. The work demonstrates improved management but does not emphasize edge deployment or multi-domain telemetry for fine-grained predictions. Likewise, [ref4] proposes a modeling and deployment framework for digital twins in 5G networks, creating virtual replicas for monitoring and optimization. Although applicable to 6G, it focuses on physical layer aspects without addressing service-level slicing dynamics. Our DTaaS extends DTs to a modular, edge-embedded service, incorporating cross-layer telemetry for slice-specific predictive orchestration.
@@
[ref1] RL-based dynamic resource allocation for slicing Reactive, may cause SLA violations in dynamic 6G
[ref2] Hybrid DL for slicing with security and latency focus No integration with real-time digital twins
[ref3] DT-driven orchestration for end-to-end slicing Lacks edge deployment and multi-domain telemetry
[ref4] DT modeling for 5G virtual replicas Focus on physical layer, not slice-level orchestration
[ref5] Deep RL for online slice resource allocation Centralized, unsuitable for URLLC latency
[ref6] Deep RL for flexible beyond-5G slicing Static training, lacks continuous feedback
[ref7] Federated learning for slicing in vehicular networks No synchronized twin representation for full state tracking
[ref8] Zero-touch orchestration for large-scale slices Focus on policies, not predictive SLA management
Predictive orchestration is essential for adaptive 6G networks. In [ref5], deep reinforcement learning is applied for online resource allocation in network slicing, achieving efficient adaptation to varying demands. Its centralized nature, however, introduces latency issues unsuitable for URLLC slices. In [ref6], deep RL is used for network slicing in beyond-5G and 6G systems to flexibly accommodate wireless services. While scalable, it relies on static training without continuous feedback loops for SLA risk prediction. Our framework deploys lightweight SDTs at the edge, merging sequential models with closed-loop control to reduce latency and bolster SLA compliance.
AI-native paradigms for 6G orchestration have also gained attention. In [ref7], federated learning is employed for network slicing in 6G systems tailored to autonomous vehicles, improving connectivity through distributed training. The approach enhances privacy but lacks synchronized twin representations for holistic state tracking. Similarly, [ref8] proposes a framework for zero-touch management and orchestration of massive network slice deployments in 6G, supporting automated lifecycle handling. It prioritizes policy enforcement over predictive resource management. Our DTaaS distinguishes itself by integrating SDTs into the orchestration loop, facilitating proactive scaling and reconfiguration via real-time telemetry and analytics.
@@ V-A Simulation Environment
To evaluate the proposed DTaaS framework, a simulation environment was developed in Python using the SimPy discrete-event simulation library [xx9617810]. The testbed emulates a 6G slicing scenario with heterogeneous services, including eMBB, URLLC, and mMTC. Each service category is instantiated as an independent network slice with unique traffic characteristics, latency constraints, and reliability targets.
@@ V-A1 Baseline Methods
  • Centralized Deep Reinforcement Learning (C-DRL): Inspired by recent studies such as [ref5, ref6], this baseline applies a centralized deep reinforcement learning model trained to allocate resources among active slices. It learns optimal allocation policies over time but lacks real-time predictive awareness and introduces control latency due to centralized inference. Full RL implementations in live networks are complex; therefore, this baseline simulates the logical behavior of centralized agents rather than training on live data, enabling fair comparison with DTaaS.